# ClimaQA: An Automated Evaluation Framework for Climate Foundation Models

### v1

@@ A.2 Training Details
We used Llama3.1-8B and Mistral-7B-v0.3 as our base models and performed continued pre-training and fine-tuning on them. We utilized the Low-Rank Adaptation (LoRA) [Hu et al., 2021] technique for efficient continued pre-training and fine-tuning.
A.2.1 Continued Pre-training on Graduate Textbook Data
@@
A.2.2 Fine-tuning on ClimaQA-Silver
@@
A.3 ClimaGen Prompts
A.3.1 QA Generation
@@
4. Do not include phrases like "According to the provided context…"
@@ 5. Do not make any references to the given context in the question
8. Use "None of the Above" or "All of the Above" sparingly.
@@
- Do not include phrases such as "According to the provided context.."
@@ A.3.2 Complexity Addition
A.4 Automated Annotation Details
A.4.1 Evaluator Model
@@
A.4.2 Cloze Annotation
To automate cloze annotation, we fine-tuned gpt-4o-mini with the following prompt to pick the scientific term to be masked. The dataset consisted a total of 160 expert-labeled examples.
@@ A.4.2 Cloze Annotation
[linewidth=1pt] You are a climate cloze generator that marks a scientific term from the given scientific statement to be masked for cloze question-answering The scientific term has to be a single word from the given statement that has a significant impact if removed You will be provided the following as the input::
Statement: $\langle statement\rangle$ Respond with just one word

### v2

@@ A.3 Training Details
We used Llama3.1-8B and Mistral-7B-v0.3 as our base models and performed continued pre-training and fine-tuning on them. We utilized the Low-Rank Adaptation (LoRA) (Hu et al., 2021) technique for efficient continued pre-training and fine-tuning.
A.3.1 Continued Pre-training on Graduate Textbook Data
@@
A.3.2 Fine-tuning on ClimaQA-Silver
@@
A.4 ClimaGen Prompts
A.4.1 QA Generation
@@
4. Do not include phrases like ”According to the provided context…”
@@ 5. Do not make any references to the given context in the question
8. Use ”None of the Above” or ”All of the Above” sparingly.
@@
- Do not include phrases such as ”According to the provided context..”
@@ A.4.2 Complexity Addition
A.5 Annotation
A.5.1 Expert Validation
During the validation process of the ClimaQA-Gold dataset for both multiple-choice and freeform questions, experts were asked to select predefined reasons for rejecting a generated QA pair. They also had the option to provide a custom reason for rejection. The figure below illustrates the various types of errors made by the ClimaGen pipeline during the generation phase. While most rejection reasons highlight limitations of the generator LLM, the category of bad context points to instances where the seed context used for QA generation was inherently flawed or lacked meaningful information. Addressing this issue through improved preprocessing techniques is a key area for future work.
Figure 7: The distribution of different reasons selected for scientific invalidity of generated QA pairs by domain experts.
A.5.2 Automated Annotation - MCQ and Freeform
@@
A.5.3 Automated Annotation - Cloze
To automate cloze annotation, we fine-tuned gpt-4o-mini with the following prompt to pick the scientific term to be masked. The dataset consisted of a total of 160 expert-labeled examples.
@@ A.5.3 Automated Annotation - Cloze
[linewidth=1pt] You are a climate cloze generator that marks a scientific term from the given scientific statement to be masked for cloze question-answering The scientific term has to be a single word from the given statement that has a significant impact if removed You will be provided the following as the input:
Statement: $\langle statement\rangle$
Respond with just one word